OpenStudy (anonymous):

If one coin is flipped and then a second coin is flipped, what is the probability that both coins turn up tails?

OpenStudy (alex6799):

25% if it was one coin then it would be 50/50 but since there are 2 coins you split the possibility again so 1/4=25%

OpenStudy (anonymous):

probability is = posibility of that event/ total no of events =1/4=0.25